# ANAH-v2: Scaling Analytical Hallucination Annotation of Large Language Models

### v1

@@ Abstract
Large language models (LLMs) exhibit hallucinations in long-form question-answering tasks across various domains and wide applications. Current hallucination detection and mitigation datasets are limited in domains and sizes, which struggle to scale due to prohibitive labor costs and insufficient reliability of existing hallucination annotators. To facilitate the scalable oversight of LLM hallucinations, this paper introduces an iterative self-training framework that simultaneously and progressively scales up the hallucination annotation dataset and improves the accuracy of the hallucination annotator. Based on the Expectation Maximization (EM) algorithm, in each iteration, the framework first applies a hallucination annotation pipeline to annotate a scaled dataset and then trains a more accurate hallucination annotator on the dataset. This new hallucination annotator is adopted in the hallucination annotation pipeline used for the next iteration. Extensive experimental results demonstrate that the finally obtained hallucination annotator with only 7B parameters surpasses the performance of GPT-4 and obtains new state-of-the-art hallucination detection results on HaluEval and HalluQA by zero-shot inference. Such an annotator can not only evaluate the hallucination levels of various LLMs on the large-scale dataset but also help to mitigate the hallucination of LLMs generations, with the Natural Language Inference (NLI) metric increasing from 25% to 37% on HaluEval.  [^1] Dataset, code, and model are released at https://github.com/open-compass/ANAH.
@@ 1 Introduction
Extensive experimental results show that our enhanced annotator significantly outperforms existing models, including the advanced GPT-4, in terms of accuracy. Our annotator not only performs best on the in-domain fine-grained hallucination annotation dataset ANAH (89.24%) but also obtains new state-of-the-art (SOTA) results on HaluEval (81.54%) and HalluQA (94.44%) under zero-shot setting. In addition, the annotator automates the hallucination evaluation on the dataset, offering a comprehensive benchmark for the research community to evaluate the hallucination levels of numerous open-source models, providing a practical reference for future hallucination mitigation of LLMs. With a simple reranking strategy with the annotator, we reduce the hallucination of the final LLM generations from 25% to 37%.
@@ 2 Related Work
Self-improvement of Large Models. As Large Language Models (LLMs) become more and more powerful, the community starts to explore different strategies to achieve the self-improvement of LLMs, i.e., to improve the LLMs using the supervision from LLMs. For example, existing works have explored self-alignment using LLMs with ethical principles [63, 3, 52]. There are also methods [22, 50, 64, 34] strengthen LLM’s capabilities on tasks such as reasoning by training the LLMs on the high-quality responses from themselves on the same questions. In the field of computer vision, SAM [32] introduces manual and model-assisted labeling to expand the image segmentation dataset and enhance the performance of image segmentation models. However, the application of self-improvement is under-explored in fine-grained hallucination annotation. This field is challenging for automatic annotators due to its meticulous nature, which requires fine validation with long documents. It is also noteworthy that most self-improvement works require extra resources such as human labor or a supplementary model [32, 34]. In contrast, our pipeline is self-sufficient, relying solely on the annotator model and the initial dataset.
@@ 3.2 Expectation-Maximization Algorithm
M-Step. Following the robust estimation in the E-step, the M-step updates the model parameters to maximize the likelihood of the selected output $y^{*}$. Combining Eq. 1 and Eq 2, we formulate the parameter update strategy at iteration $t$:
@@ 4.2 Overall Results
The last 3 rows of Tab. 2 illustrate the performance of ANAH-v2 at each stage of Data Scaling in § 3.3. The performance progressively improves with the increasing dataset number (see in Tab. 1) in successive stages. This trend underscores the scalability and effectiveness of our hallucination annotation framework. Remarkably, ANAH-v2 surpasses GPT-4 with the F1 of 87.78% and the accuracy of 88.03% at Stage 2  [^4] Notably, GPT-4 annotations are used as pre-annotation in ANAH, which makes the RougeL and BERTScore higher than the zero-shot setting. We use “accuracy” as our primary metric because the type judgment is determinative of the annotation quality. For example, an annotation that wrongly judges type (low accuracy) but finds the correct reference fragment (high RougeL/BERTScore) remains completely unacceptable. . Eventually, we achieve the F1 of 89.30% and the accuracy of 89.55% at Stage 3.
@@ 4.5 Application
We find that all models exhibit superior performance in English compared to Chinese, underscoring the need for further research to understand and mitigate language-dependent discrepancy. The performances of all models with reference documents are better than those without. Qwen1.5-14B achieves the lowest hallucination rate when using reference documents (5.33%) and Deepseek-67B achieves the lowest hallucination rate when reference documents are not provided (47.17%). Moreover, we find no clear trend in the performance distribution across four categories of topics. In addition, the results of different stages of annotators in Tab. A1,  A2, and  7 show that there is a consistent trend and fixed biased ordering relationship between LLMs, thus confirming the reliability of our assessment method. More details are in Appendix D.
@@ Appendix C Implementation Details
Appendix D Hallucination Evaluation
To assess the reliability of our hallucination annotator, we measure the hallucination levels of the above LLMs using the annotator from different stages. Tab. A1,  A2, and  7 show the results measured by annotator ANAH-v2 from Stage 1, 2, and 3, respectively. The trends in these three tables are consistent where Qwen1.5-14B achieves the lowest hallucination rate with reference documents and DeepseekLM-67B achieves the lowest hallucination rate without reference documents. This consistency and fixed biased ordering relationship between LLMs confirm the reliability of our assessment method.
Appendix E Limitation
@@
Appendix F Broader Impacts

### v2

@@ Abstract
Large language models (LLMs) exhibit hallucinations in long-form question-answering tasks across various domains and wide applications. Current hallucination detection and mitigation datasets are limited in domain and size, which struggle to scale due to prohibitive labor costs and insufficient reliability of existing hallucination annotators. To facilitate the scalable oversight of LLM hallucinations, this paper introduces an iterative self-training framework that simultaneously and progressively scales up the annotation dataset and improves the accuracy of the annotator. Based on the Expectation Maximization algorithm, in each iteration, the framework first applies an automatic hallucination annotation pipeline for a scaled dataset and then trains a more accurate annotator on the dataset. This new annotator is adopted in the annotation pipeline for the next iteration. Extensive experimental results demonstrate that the finally obtained hallucination annotator with only 7B parameters surpasses GPT-4 and obtains new state-of-the-art hallucination detection results on HaluEval and HalluQA by zero-shot inference. Such an annotator can not only evaluate the hallucination levels of various LLMs on the large-scale dataset but also help to mitigate the hallucination of LLMs generations, with the Natural Language Inference metric increasing from 25% to 37% on HaluEval.  [^1] Dataset, code, and model are released at https://github.com/open-compass/ANAH.
@@ 1 Introduction
Extensive experimental results show that our enhanced annotator significantly outperforms existing models, including the advanced GPT-4, in terms of accuracy. Our annotator not only performs best on the in-domain fine-grained hallucination annotation dataset ANAH (89.24%) but also obtains new state-of-the-art (SOTA) results on HaluEval (81.54%) and HalluQA (94.44%) under zero-shot setting. In addition, the annotator automates the hallucination evaluation on the dataset, offering a comprehensive benchmark for the research community to evaluate the hallucination levels of numerous open-source models, providing a practical reference for future hallucination mitigation of LLMs. Using a simple reranking strategy with the annotator, we reduce the hallucination of the final LLM generations on HaluEval, with the NLI metric increasing from 25% to 37%.
@@ 2 Related Work
Self-improvement of Large Models. As Large Language Models (LLMs) become more and more powerful, the community starts to explore different strategies to achieve the self-improvement of LLMs, i.e., to improve the LLMs using the supervision from LLMs [12, 24, 28, 53, 50]. For example, existing works have explored self-alignment using LLMs with ethical principles [69, 3, 58]. There are also methods [25, 56, 70, 38, 13] strengthen LLM’s capabilities on tasks such as reasoning by training the LLMs on the high-quality responses from themselves on the same questions. In the field of computer vision, SAM [36] introduces manual and model-assisted labeling to expand the image segmentation dataset and enhance the performance of image segmentation models. However, the application of self-improvement is under-explored in fine-grained hallucination annotation. This field is challenging for automatic annotators due to its meticulous nature, which requires fine validation with long documents. It is also noteworthy that most self-improvement works require extra resources such as human labor or a supplementary model [36, 38]. In contrast, our pipeline is self-sufficient, relying solely on the annotator model and the initial dataset.
@@ 3.2 Expectation-Maximization Algorithm
M-Step. Following the robust estimation in the E-step, the M-step updates the model parameters to maximize the likelihood of the selected output $y^{*}$. Combining Eq. 1 and Eq. 2, we formulate the parameter update strategy at iteration $t$:
@@ 4.2 Overall Results
The last 3 rows of Tab. 2 illustrate the performance of ANAH-v2 at each stage of Data Scaling in § 3.3. The performance progressively improves with the increasing dataset number (see in Tab. 1) in successive stages. This trend underscores the scalability and effectiveness of our hallucination annotation framework. Remarkably, ANAH-v2 surpasses GPT-4 with the F1 of 87.78% and the accuracy of 88.03% at Stage 2. Eventually, we achieve the F1 of 89.30% and the accuracy of 89.55% at Stage 3.
Notably, the RougeL and BERTScore of GPT-4 are higher than ANAH-v2. Because GPT-4 is used for the initial pre-annotation during the construction of ANAH [29]. Subsequently, humans refine these pre-annotations, and humans tend to not change the pre-annotations. This methodology inherently aligns the final ’golden’ answers closely with the outputs by GPT-4. Therefore, we tend to use “accuracy” as our primary metric because the type judgment is determinative of the annotation quality. For example, an annotation that wrongly judges type (low accuracy) but finds the correct reference fragment (high RougeL/BERTScore) remains completely unacceptable. In addition, we conduct an LLM-based evaluation to exclude the similarity due to pre-annotations in Appendix D.
@@ 4.5 Application
We find that all models exhibit superior performance in English compared to Chinese, underscoring the need for further research to understand and mitigate language-dependent discrepancy. The performances of all models with reference documents are better than those without. Qwen1.5-14B achieves the lowest hallucination rate when using reference documents (5.33%) and Deepseek-67B achieves the lowest hallucination rate when reference documents are not provided (47.17%). Moreover, we find no clear trend in the performance distribution across four categories of topics. In addition, the results of different stages of annotators in Tab. A2,  A3, and  7 show that there is a consistent trend and fixed biased ordering relationship between LLMs, thus confirming the reliability of our assessment method. More details are in Appendix E.
@@ 5 Conclusion and Future Work
Acknowledgement
We thank the anonymous reviewers and area chair for their helpful comments. This project is funded in part by the Hong Kong Generative AI Research and Development Center (HKGAI) under the Innovation and Technology Commission (ITC)’s InnoHK. Dahua Lin is a PI of HKGAI under the InnoHK. This project is also supported by the Shanghai Artificial Intelligence Laboratory. The authors would like to thank Zehui Chen and Kuikun Liu for their valuable suggestions and comments.
@@ Appendix C Implementation Details
Appendix D Addtional Evaluation for the Overall Results
Model Score
GPT-4 84.39
ANAH-7B 80.60
ANAH-20B 81.51
ANAH-V2-Stage1 83.63
ANAH-V2-Stage2 84.54
ANAH-V2-Stage3 86.36
Table A1: The score assessing the consistency of generated reference points with the source documents.
To exclude the similarity due to the impact of pre-annotations described in § 4.2, we conduct an LLM-based evaluation to assess the consistency of generated reference points with the source document. Specifically, we followed the prompt in FactScore [46], which aims to clarify whether the generated reference points are supported by the given source document, rather than simply calculating the similarity between them using metrics such as RougeL or BERTScore. We employ the InternLM2-7B-Chat [7] as the estimator.
As shown in Tab. A1, the results indicate that the reliability of our model’s generated reference points progressively improves and ultimately exceeds that of GPT4. This trend is consistent with F1 and ACC in Tab 2.
Appendix E Hallucination Evaluation
To assess the reliability of our hallucination annotator, we measure the hallucination levels of the above LLMs using the annotator from different stages. Tab. A2,  A3, and  7 show the results measured by annotator ANAH-v2 from Stage 1, 2, and 3, respectively. The trends in these three tables are consistent where Qwen1.5-14B achieves the lowest hallucination rate with reference documents and DeepseekLM-67B achieves the lowest hallucination rate without reference documents. This consistency and fixed biased ordering relationship between LLMs confirm the reliability of our assessment method.
Appendix F Limitation
@@
In addition, the EM algorithm, which is the theoretical foundation of our framework, may also introduce some problems. For example, the EM algorithm is sensitive to initial conditions, which would impact the convergence process. Although we employ many methods to ensure the stability of the training process, such as selecting a high-quality, human-labeled hallucination dataset as the seed and using a progressive scaling strategy, we cannot claim that we have eventually converged to a globally optimal solution. Moreover, the iterative EM algorithm requires computational effort. Our method uses 32 A100 GPUs to iteratively train the 7B model. It took approximately 100 hours for inference and training. Based on the price of the computing platform Lambda (1.29 USD per GPU per hour), it costs 4,128 USD. However, using the "manual + GPT4-assisted" annotation model, as described in ANAH [29] (0.9 USD and 20 minutes per annotation), it would take 177,237 USD and 65,643 hours to reach the size of the dataset in our work. So we believe our method is a better trade-off between computing resources and labour+API costs, which is acceptable.
Appendix G Broader Impacts